OpenStudy (anonymous):

Functions f(x) and g(x) are shown below. f(x) = x^2 g(x) = x^2 + 8x + 16 In which direction and by how many units should f(x) be shifted to obtain g(x)? Left by 4 units Right by 4 units Left by 8 units Right by 8 units

OpenStudy (anonymous):

@TwinYang @Cosmichaotic

OpenStudy (anonymous):

\[g(x) = x^2 - 8x + 16\] You can see that the above function is a perfect square. Hope you remember that \[(a-b)^2=a^2-2ab+b^2\] x^2 - 8x + 16 is the same with a=x and b=4 \[x^2 - 8x + 16=(x-4)^2\] So the shift from x^2 is by 4 units.

OpenStudy (anonymous):

So now how do I determine if it is a right shift or a left shift? Are those just other ways of saying negative or positive?

OpenStudy (anonymous):

@TwinYang ?

OpenStudy (anonymous):

left

OpenStudy (anonymous):

Okay thanks!